Some students planned a picnic the total budjet for hiring a bus was 1440 Later on  8 of these refused to go and instead paid their total share of money towards the fee of one economically weaker student of their and thus the cost of each member who went for picnic increased by 30.
a)How many students attended the picnic?
b)How much money in total was paid towards the fees?which value is reflected in this question?

Grade:10

1 Answers

Arun
4 years ago
Dear student
 

total- 1440

no of students who where supposed to attend - x

no of students who attended-x-8

fee paid by each member in case 1- 1440/x

fee paid by each member in case2-1440/x-8

therefore acc. to the problem

1440/x-8=(1440/x)+30

= (1440/x-8) -1440/x=30

x^2-8x-384 (we will get this equation in the end)

after factorizing u will get

x=24 x=-16 .... therefor x=24

(1) no of stud. who attended = 24-8=16

(2) total fee by each member=1440/16=180 rs